How many characters per second can be transmitted over a 1200-baud line in each of the following modes? (Assume a character code of eight bits.)

a. Synchronous serial transmission.

b. Asynchronous serial transmission with two stop bits.

c. Asynchronous serial transmission with one stop bit.



1
Expert's answer
2021-11-24T00:36:02-0500

a. Synchronous serial transmission. Also draw timing diagram.

Baud Rate = 9600 bps = 9600 bits per second = 1200 bytes per second

This means a maximum of 1200 characters (ASCII codes ) can be transmitted in one second at the baud rate = 9600 bps.

Data transfer is initiated by by MASTER to SLAVE over a synchronus clock pulse

Number of 8-bit characters that can be transmitted per second = (9600)/8 =

1200 character bytes.






b. Asynchronous serial transmission with two stop bits. Draw timing diagram

Total Data To send in 1 frame = 1 bit(start) + 8 bits(char size) + 1 bit(Parity) + 2 bits (Stop) = 12 bits.

 

Number of 8-bit characters that can be transmitted per second = (9600)/12 =

800 bytes.



c. Asynchronous serial transmission with one stop bit. Draw timing diagram.

Total Data To send in 1 frame = 1 bit(start) + 8 bits(char size) + 1 bit(Parity) + 1 bits (Stop) = 11 bits

Number of 8-bit characters that can be transmitted per second = (9600)/11 =

872 character bytes.